# Doublons Bloch oscillations in the mass-imbalanced extended Fermi-Hubbard model

### v1

@@ Abstract
Interactions between particles normally induce the decay of the particles Bloch oscillations (BOs) in a periodic lattice. In the large on-site interactions region, spin-$1/2$ fermions may form into doublon bound state and undergoes doublon BOs in the present of tilted potential. Here we investigate the impact of nearest-neighbor interaction $V$ on the multi-doublon BOs in a mass-imbalanced extended Fermi-Hubbard model. We derive a doublons effective Hamiltonian, and show that a minor change of $V$ can alter the dynamic behaviors of doublons qualitatively. Notably, at a resonance point, the doublons act like free hard-core bosons. Under a tilted potential, the system may exhibit different types of multi-doublon BOs at or deviation from the resonance point with different initial states. Numerical results are presented to demonstrate our conclusions in the one- and two-dimensional systems.
@@ I Introduction
Bloch oscillations (BOs) of electrons in crystal lattices were first predicted by Felix Bloch in 1929 [1]. When a finite dc electric field is applied to the ideal crystal lattice, the eigenstates of electrons become localized, accompanied by the ladder-like energy levels, i.e., the Wannier-Stark (WS) ladder [2], resulting in the coherent motion of wave packet, which is the underlying mechanism of the BOs. Until sixty years after being predicted, the BOs were experimentally observed in semiconductor superlattices [3, 4] and optical lattices [5, 6, 7, 8]. In recent years, the BOs have been intensively explored in strongly correlated systems, including quantum spin chain [9, 10, 11, 12, 13, 14, 15, 16], interacting boson and fermion systems [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32], in both theoretical and experimental aspects.
@@
In general, interactions between particles induce the decoherence of BOs [17, 18, 19, 20, 21, 22]. Nevertheless, a strong interaction strength may bound two particles into a doublon [33, 34, 35, 29, 30, 36, 37, 38, 39, 40, 41, 42] or a bound pair [43, 44, 45, 46, 27, 47, 48, 49] as high-energy excitations, in the boson and fermion systems, respectively. The BOs of bound particles in these systems exhibit characteristics different from the free-particle situations, for example, the fractional BOs [27, 50] or frequency doubling phenomena [23, 24, 25, 51] have been observed. The Fermi-Hubbard model [52, 53] describes a typical strongly correlated electronic system, with hopping amplitude $J$ and on-site interaction with strength $U$. A more general consideration leads to the extended Fermi-Hubbard model [29, 36, 54, 55, 56, 57, 58], which includes the nearest-neighbor (NN) interaction with strength $V$. Obviously, both the on-site and NN interactions can affect the behavior of many-body BOs. The aforementioned research works have studied the BOs in the (extended) Fermi-Hubbard chains, but mainly focused on the one doublon or two fermions case. It is natural and interesting to take the doublon-doublon interaction into account in the dynamics of doublons BOs.
@@
This paper aims to investigate the doublons BOs in the extended Hubbard model with imbalanced mass of spin-up and spin-down fermions [37, 59, 60, 61, 62, 63, 64, 65, 66, 67], under a tilted potential. We derive a doublons effective Hamiltonian in the strong on-site interaction situation to capture the physics of the doublons dynamics. We show that the NN interaction induces the bound-doublon eigenstates, which dominate the doublons dynamics. It is observed that (i) at a resonance point, the bound-doublon eigenstates disappear, and the doublons act like free hard-core bosons; (ii) when the NN interaction is strong enough, the energy levels of bound doublons are completely isolated from the scattering band of doublons. These properties enable the doublons BOs to occur after applying the tilted potential for the systems in the above two cases. We present the numerical simulations of time evolutions to verify our analysis. For the above two cases, the breathing modes of BOs exhibit the same pattern of doublon density, but with significant difference of mean-square displacements (MSD) of doublons in one dimension ($1$D). The numerical results of doublons BOs for the system on a square lattice are also presented and discussed. Our results show that the multi-doublon BOs can survive in spite of the present of the NN interaction of fermions.
@@ II Model and doublons effective Hamiltonian
is the extended Fermi-Hubbard model with imbalance mass. Figure 1 presents a schematic illustration of the system in $1$D. For a $1$D system, the dc field term with strength $F$ takes the form of tilted potential
@@
which is readily feasible in the experiments of ultracold atoms [68, 22, 69]. Here, $\hat{c}_{i,\sigma}^{\dagger}$ ($\hat{c}_{i,\sigma}$) is the creation (annihilation) operator of fermion with spin $\sigma$ on the site with index $i$; the summation $\sum_{\left\langle i,j\right\rangle}$ is over the NN sites with indices $i$ and $j$; and $\hat{n}_{i,\sigma}=\hat{c}_{i,\sigma}^{\dagger}\hat{c}_{i,\sigma}$, $\hat{n}_{i}=\hat{n}_{i,\uparrow}+\hat{n}_{i,\downarrow}$ are the corresponding fermion number operators. The strength of on-site interaction and NN interaction are $U$ and $V$, respectively, which can be positive (repulsion) or negative (attraction). The mass of the spin-up and spin-down fermions are different, manifesting as the difference of the corresponding hopping amplitudes. This has been realized with Fermi gas of ${}^{171}\mathrm{Yb}$ atoms in a state-dependent optical lattice recently [65]. We assume $J_{\downarrow}>J_{\uparrow}>0$ in the following discussions. It has been shown that the heavy particles acting as an effective disorder potential for the light ones in a extremely strong mass imbalance [59]. The system conserves the total fermion number $\hat{n}=\sum_{i}\hat{n}_{i}$, as well as the total spin polarization $\hat{s}^{z}=\sum_{i}(\hat{n}_{i,\uparrow}-\hat{n}_{i,\downarrow})/2$.
@@
The standard Hubbard model in 1D was exactly solved by Lieb and Wu with the Bethe ansatz method [70]. The imbalance of mass $J_{\uparrow}\neq J_{\downarrow}$ and nonzero NN interaction $V\neq 0$ resists the exact analytical treatments so far. Nevertheless, we pursue the perturbation and numerical solutions. In this work, we focus on the dynamics of the system with multidoublon, i.e., each site of the system is either doubly occupied or empty. This invariant subspace, referred to as doublons subspace, is approximately achieved when considering the strong on-site interaction region $\left|U\right|\gg J_{\sigma},\left|V\right|,\left|F\right|$. In the following, we derive the effective Hamiltonian of this case.
To explore the feasibility of the doublons BOs in the present model, we shall derive the doublons effective Hamiltonian. We firstly focus on the Hamiltonian without the tilted potential. In the strong on-site interaction region, the states in the doublons subspace are well separated from others in the spectrum. The doublons effective Hamiltonian for $\hat{H}$ can be obtained perturbatively through the Schrieffer-Wolff transformation [71, 72, 73, 74, 75]
@@
where $\mathrm{i}=\sqrt{-1}$. The terms that change the number of doublons can be eliminated by properly selecting the Hermitian operator $\hat{S}$. Up to second order, the doublons effective Hamiltonian reads (see the APPENDIX for detailed calculation)
@@
where $d_{i}^{\dagger}=\hat{c}_{i,\uparrow}^{\dagger}\hat{c}_{i,\downarrow}^{\dagger}$ is the creation operator of doublon; the effective hopping amplitude and NN interaction of doublons are $J_{\mathrm{eff}}=2J_{\uparrow}J_{\downarrow}/U$ and $V_{\mathrm{eff}}=V-(J_{\uparrow}^{2}+J_{\downarrow}^{2})/U$, respectively. This means that the doublons move in the lattice through the second-order processes, and all the three parameters of the Hubbard model play a crucial part in the doublon-doublon interaction. Clearly, the doublons satisfies the statistics of hard-core bosons, that is
@@ III.1 Two-doublon Bloch oscillations
$+\frac{1}{2}V_{\mathrm{eff}}\left|\phi_{1}(K)\right\rangle\left\langle\phi_{1}(K)\right|,$
@@
$\left|\phi_{r}(K)\right\rangle=\frac{1}{\sqrt{L}}e^{\mathrm{i}Kr/2}\sum_{j}e^{\mathrm{i}Kj}\hat{d}_{j}^{\dagger}\hat{d}_{j+r}^{\dagger}\left|\text{{Vac}}\right\rangle,$ (8)
where the vacuum state is defined as $\hat{d}_{i}\left|\text{{Vac}}\right\rangle=0$, and $r=1,2,3,...,L-1$ is the distance between two doublons. Note that $\hat{H}_{\mathrm{eff}}^{(2)}(K)$ can be regarded as the single-particle Hamiltonian of a tight-binding chain with hopping amplitude $2J_{\mathrm{eff}}\cos(K/2)$, and with impurity potential $V_{\mathrm{eff}}/2$ at one end. Such impurity model supports bound state in certain parameter region. In order to solve the eigenvalue equation $\hat{H}_{\mathrm{eff}}^{(2)}(K)\left|\psi(K)\right\rangle=E(K)\left|\psi(K)\right\rangle$, we expand the eigenstate $\left|\psi(K)\right\rangle$ as
@@
$+\left[\frac{1}{2}V_{\mathrm{eff}}\delta_{r,1}-E(K)\right]C_{r}(K)=0,$ (10)
@@
$\beta(K)=\ln\left[\frac{V_{\mathrm{eff}}}{4J_{\mathrm{eff}}\cos\left(K/2\right)}\right],$ (11)
@@
$E_{\textrm{bou.}}(K)=\frac{1}{2}V_{\mathrm{eff}}+\frac{8J_{\mathrm{eff}}^{2}\cos^{2}\left(K/2\right)}{V_{\mathrm{eff}}}.$ (12)
The bound-state condition $\beta(K)>0$ requires $\left|V_{\mathrm{eff}}\right|>4\left|J_{\mathrm{eff}}\cos\left(K/2\right)\right|$. In this case, the doublons bound states are induced by the NN interaction $V$. In Fig. 2, we present the two-doublon spectrum for various NN interaction strengths, which can be obtained by diagonalizing the two-doublon effective Hamiltonian Eq. (7) or solving Eq. (10). In Fig. 2(a), we can see that when $V_{\mathrm{eff}}=0$, the spectrum is all composed of scattered band of doublons. Increasing $\left|V\right|$ would induce the isolated energy levels, which are the energies of bound doublons [see Fig. 2(b)]. When $\left|V_{\mathrm{eff}}\right|>4\left|J_{\mathrm{eff}}\cos\left(K/2\right)\right|$ for all $K$, the bound-doublon band is completely isolated from the scattered band [see Fig. 2(c)]. In the following, we shall investigate the BOs in the parameter regions of Figs. 2(c) and (a) sequentially.
Figure 2: Two-doublon spectrum as a function of momentum $K$ for various NN interactions, which are obtained from the exact diagonalization of the two-doublon effective Hamiltonian in Eq. (7). Parameters of the systems are (a) $V=-1/6$, (b) $V=-0.8$ and (c) $V=-2$. Other parameters are $L=100$, $U=-6$, $J_{\downarrow}=1$ and $J_{\uparrow}=0.5$. The doublons scattered bands and bound-doublon energy levels are indicated by the black and red lines, respectively. Here the energy is in unit of $J_{\downarrow}$.
Here we first focus on the dynamics of the two bound doublons. For a large NN interaction $V$, we have $\left|V_{\mathrm{eff}}\right|\gg 4\left|J_{\mathrm{eff}}\cos\left(K/2\right)\right|$. In this case, two doublons in the bound band are strongly bound. Then we are able to write down the two-doublon effective Hamiltonian under the strongly bound basis $\left|\varphi_{j}\right\rangle=\hat{d}_{j}^{\dagger}\hat{d}_{j+1}^{\dagger}\left|\text{{Vac}}\right\rangle$. We divide the effective Hamiltonian Eq. (5) into two parts, i.e., $\hat{H}_{\mathrm{eff}}^{(0)}=V_{\mathrm{eff}}\hat{n}_{i}\hat{n}_{j}/2$ and $\hat{H}_{\mathrm{eff}}^{\prime}=J_{\mathrm{eff}}(\hat{d}_{i}^{\dagger}\hat{d}_{j}+\hat{d}_{j}^{\dagger}\hat{d}_{i})$, and treat $\hat{H}_{\mathrm{eff}}^{\prime}$ as a perturbation. According to the degenerate perturbation theory, the effective Hamiltonian is given by the following formula
@@
up to second order $J_{\mathrm{eff}}^{2}/V_{\mathrm{eff}}$. Here the projectors are defined as $\hat{P}=\sum_{j}\left|\varphi_{j}\right\rangle\left\langle\varphi_{j}\right|$ and $\hat{Q}=\hat{1}-\hat{P}$, and the energy of the unperturbed state is $E_{\mathrm{eff}}^{(0)}=V_{\mathrm{eff}}/2$. Direct calculation shows that
$\hat{H}_{\mathrm{eff}}^{(2,\mathrm{bou.})}\approx\frac{2J_{\mathrm{eff}}^{2}}{V_{\mathrm{eff}}}\sum_{j}\left(\left|\varphi_{j}\right\rangle\left\langle\varphi_{j+1}\right|+\left|\varphi_{j+1}\right\rangle\left\langle\varphi_{j}\right|\right),$ (14)
in which the constant potential that is proportional to $\sum_{j}\left|\varphi_{j}\right\rangle\left\langle\varphi_{j}\right|$ is omitted. Now the bound-doublon physics is clear. From Eq. (14), we can see that the bound-doublon hops in the lattice through a second-order process of the single-doublon hopping.
Figure 3: Numerical results of the time evolution of doublon density defined in Eq. (17) for different initial states with (a) delta distribution in Eq. (15) and (b) Gaussian distribution in Eq. (16). Parameters of the system are set as the same as that in Fig. 2 (c), except that the system size is $L=30$. The strength of tilted potential is taken as $F=0.001$. Parameters of the Gaussian wave packet are $\alpha=0.4$, $j_{0}=15$ and $K_{0}=\pi/2$. The time unit is $J_{\downarrow}^{-1}$, and we set $J_{\downarrow}=1$ throughout the numerical simulations in this work.
@@
$\left|\Psi_{\mathrm{a}}\left(0\right)\right>=\hat{d}_{15}^{\dagger}\hat{d}_{16}^{\dagger}\left|\text{{Vac}}\right\rangle,$ (15)
@@
The numerical results are presented in Fig. 3, which are consistent with our analysis. Figures. 3(a) and (b) show that two different initial excitations undergo the breathing mode and oscillating mode, respectively. The oscillating periods are both $T=2\pi/(4F)\approx 1570.80$.
Figure 4: Numerical results of the time evolution of doublon density for different NN interaction $V$, tilted potential $F$ and doublon number. The system sizes for the two-doublon cases in (a)-(d) and four-doublon cases in (e)-(f) are taken as $L=30$ and $L=20$, respectively. Other system parameters are $U=-10$, $J_{\downarrow}=1$ and $J_{\uparrow}=0.1$.
Figure 5: Mean-square displacement between the doublons [see Eq. (18)] as a function of time for the cases in Fig. 3(a) and Figs. 4(a)-(d). The time scales are normalized by the final time of the data in each case.

### v3

# Two-doublon Bloch oscillations in the mass-imbalanced extended Fermi-Hubbard model

@@ Abstract
Interactions between particles normally induce the decay of the particles Bloch oscillations (BOs) in a periodic lattice. In the limit of strong on-site interactions, spin-$1/2$ fermions may form doublon bound states and undergo BOs in the presence of a tilted potential. Here we investigate the impact of nearest-neighbor interaction $V$ on the multi-doublon BOs in a mass-imbalanced extended Fermi-Hubbard model. We derive an effective Hamiltonian for doublons, and show that a slight change in $V$ can qualitatively alter their dynamic behaviors. Notably, at a resonance point, the doublons behave like free hard-core bosons. Under a tilted potential, the system may exhibit different types of multi-doublon BOs at or deviation from the resonance point. Numerical results are presented to demonstrate our conclusions in both one- and two-dimensional systems.
@@ I Introduction
Bloch oscillations (BOs) of electrons in crystal lattices were first predicted by Bloch in 1929 [1]. When a finite dc electric field is applied to the ideal crystal lattice, the eigenstates of electrons become localized, accompanied by the ladder-like energy levels, i.e., the Wannier-Stark (WS) ladder [2], resulting in the coherent motion of wave packet, which is the underlying mechanism of the BOs. Sixty years after being predicted, the BOs were experimentally observed in semiconductor superlattices [3, 4] and optical lattices [5, 6, 7, 8]. In recent years, the BOs have been intensively explored in strongly correlated systems, including quantum spin chain [9, 10, 11, 12, 13, 14, 15, 16], interacting boson and fermion systems [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32], in both theoretical and experimental aspects.
@@
In general, interactions between particles induce the decoherence of BOs [17, 18, 19, 20, 21, 22]. Nevertheless, a strong interaction strength may bind two particles into a doublon [33, 34, 35, 29, 30, 36, 37, 38, 39, 40, 41, 42] or a bound pair [43, 44, 45, 46, 27, 47, 48, 49] as high-energy excitations, in the boson and fermion systems, respectively. The BOs of bound particles in these systems exhibit characteristics different from the free-particle situations, for example, the fractional BOs [27, 50] or frequency doubling phenomena [23, 24, 25, 51] have been observed. The Fermi-Hubbard model [52, 53] describes a typical strongly correlated electronic system, with hopping amplitude $J$ and on-site interaction with strength $U$. A more general consideration leads to the extended Fermi-Hubbard model [29, 36, 54, 55, 56, 57, 58], which includes the nearest-neighbor (NN) interaction with strength $V$. Obviously, both the on-site and NN interactions can affect the behavior of many-body BOs. The aforementioned research works have studied the BOs in the (extended) Fermi-Hubbard chains, but mainly focused on the one doublon or two fermions case. It is natural and interesting to take the doublon-doublon interaction into account in the dynamics of doublons BOs.
@@
This paper aims to investigate the doublons BOs in the extended Hubbard model with imbalanced mass of spin-up and spin-down fermions [37, 59, 60, 61, 62, 63, 64, 65, 66, 67], under a tilted potential. We derive a doublons effective Hamiltonian in the strong on-site interaction situation to capture the physics of the doublons dynamics. Then we analyze the mechanism of BOs in the two-doublon subspace. We show that the NN interaction induces the bound-doublon eigenstates, which dominate the doublons dynamics. It is observed that (i) at a resonance point, the bound-doublon eigenstates disappear, and the doublons act like free hard-core bosons; (ii) when the NN interaction is strong enough, the energy levels of bound doublons are completely isolated from the scattering band of doublons. These properties enable the doublons BOs to occur after applying the tilted potential for the systems in the above two cases. To verify the analysis, we present the numerical simulations of time evolutions for the doublons dynamics, for the attractive on-site and NN interactions, i.e., $U<0$ and $V<0$, respectively, in the extended Fermi-Hubbard model. For the above two cases, the breathing modes of BOs exhibit the same pattern of doublon density, but with significant difference of mean-square displacements (MSD) of doublons in one dimension. The numerical results of doublons BOs for the system on a square lattice are also presented and discussed. Our results show that the multi-doublon BOs can survive in spite of the presence of the NN interaction of fermions.
@@ II Model and doublons effective Hamiltonian
is the extended Fermi-Hubbard model with imbalance mass. Figure 1 presents a schematic illustration of the system in one dimension. For a one-dimensional ($1$D) system, the dc field term with strength $F$ takes the form of tilted potential
@@
which is readily feasible in the experiments of ultracold atoms [68, 22, 69]. Here, $\hat{c}_{i,\sigma}^{\dagger}$ ($\hat{c}_{i,\sigma}$) is the creation (annihilation) operator of fermion with spin $\sigma$ on the site with index $i$; the summation $\sum_{\left\langle i,j\right\rangle}$ is over the NN sites with indices $i$ and $j$; and $\hat{n}_{i,\sigma}=\hat{c}_{i,\sigma}^{\dagger}\hat{c}_{i,\sigma}$, $\hat{n}_{i}=\hat{n}_{i,\uparrow}+\hat{n}_{i,\downarrow}$ are the corresponding fermion number operators. The strength of on-site interaction and NN interaction are $U$ and $V$, respectively, which can be positive (repulsion) or negative (attraction). The mass of the spin-up and spin-down fermions are different, manifesting as the difference of the corresponding hopping amplitudes. This can be experimentally realized with Fermi gas, such as gas of fermionic atoms ${}^{171}\mathrm{Yb}$, ${}^{173}\mathrm{Yb}$, ${}^{87}\mathrm{Sr}$ or ${}^{40}\mathrm{K}$, in the state-dependent optical lattices [65, 70, 71, 72, 73]. We assume $J_{\downarrow}>J_{\uparrow}>0$ in the following discussions, and focus on the the attractive on-site and NN interactions case, i.e., $U<0$ and $V<0$, respectively. The range of on-site interaction strength $U/J_{\downarrow}\in[-20,5]$ is experimentally accessible via magnetic-field tuning of the orbital Feshbach resonance, as, e.g., in the fermionic ${}^{171}\mathrm{Yb}$ atoms system [65]. It has been shown that the heavy particles acting as an effective disorder potential for the light ones in a extremely strong mass imbalance [59]. The system conserves the total fermion number $\hat{n}=\sum_{i}\hat{n}_{i}$, as well as the total spin polarization $\hat{s}^{z}=\sum_{i}(\hat{n}_{i,\uparrow}-\hat{n}_{i,\downarrow})/2$. The imbalance of mass breaks the $\textrm{SU}(2)$ symmetry of the model, resulting in different propagation speeds of fermions with different spins in the lattice in general.
@@
The standard Hubbard model in 1D was exactly solved by Lieb and Wu with the Bethe ansatz method [74]. The imbalance of mass $J_{\uparrow}\neq J_{\downarrow}$ and nonzero NN interaction $V\neq 0$ resists the exact analytical treatments so far. Nevertheless, we pursue the perturbation and numerical solutions. In this work, we focus on the dynamics of the system with multidoublon, i.e., each site of the system is either doubly occupied or empty. This invariant subspace, referred to as doublons subspace, is approximately achieved when considering the strong on-site interaction limit $\left|U\right|\gg J_{\sigma},\left|V\right|,\left|F\right|$. In the following, we derive the effective Hamiltonian of this case.
To explore the feasibility of the doublons BOs in the present model, we derive the doublons effective Hamiltonian. We firstly focus on the Hamiltonian without the tilted potential. In the strong on-site interaction limit, the states in the doublons subspace are well separated from others in the spectrum. The doublons effective Hamiltonian for $\hat{H}$ can be obtained perturbatively through the Schrieffer-Wolff transformation [75, 76, 77, 78, 79]
@@
where $\mathrm{i}=\sqrt{-1}$. The terms that change the number of doublons can be eliminated by properly selecting the Hermitian operator $\hat{S}$. Up to second order, the doublons effective Hamiltonian reads (see APPENDIX A for detailed calculation)
@@
where $d_{i}^{\dagger}=\hat{c}_{i,\uparrow}^{\dagger}\hat{c}_{i,\downarrow}^{\dagger}$ is the creation operator of doublon; the effective hopping amplitude and NN interaction of doublons are $J_{\mathrm{eff}}=2J_{\uparrow}J_{\downarrow}/U$ and $V_{\mathrm{eff}}=V-(J_{\uparrow}^{2}+J_{\downarrow}^{2})/U$, respectively. This means that the doublons move in the lattice through the second-order processes, and all three parameters of the Hubbard model play a crucial part in the doublon-doublon interaction. Clearly, the doublons satisfies the statistics of hard-core bosons, that is
@@ III.1 Two-doublon Bloch oscillations
$+2V_{\mathrm{eff}}\left|\phi_{1}(K)\right\rangle\left\langle\phi_{1}(K)\right|,$
@@
$\left|\phi_{r}(K)\right\rangle=\frac{1}{\sqrt{L-1}}e^{\mathrm{i}Kr/2}\sum_{j}e^{\mathrm{i}Kj}\hat{d}_{j}^{\dagger}\hat{d}_{j+r}^{\dagger}\left|\text{{Vac}}\right\rangle,$ (8)
where the vacuum state is defined as $\hat{d}_{i}\left|\text{{Vac}}\right\rangle=0$, and $r=1,2,3,...,L-1$ is the distance between two doublons. Note that $\hat{H}_{\mathrm{eff}}^{(2)}(K)$ can be regarded as the single-particle Hamiltonian of a tight-binding chain with hopping amplitude $2J_{\mathrm{eff}}\cos(K/2)$, and with impurity potential $2V_{\mathrm{eff}}$ at one end. Such impurity model supports bound state in certain parameter region. In order to solve the eigenvalue equation $\hat{H}_{\mathrm{eff}}^{(2)}(K)\left|\psi(K)\right\rangle=E(K)\left|\psi(K)\right\rangle$, we expand the eigenstate $\left|\psi(K)\right\rangle$ as
@@
$+\left[2V_{\mathrm{eff}}\delta_{r,1}-E(K)\right]C_{r}(K)=0,$ (10)
@@
Figure 2: Two-doublon spectrum as a function of momentum $K$ for various NN interactions, which are obtained from the exact diagonalization of the two-doublon effective Hamiltonian in Eq. (7). Parameters of the systems are (a) $V=-5/32$, (b) $V=-0.25$ and (c) $V=-0.5$. Other parameters are $L=100$, $U=-8$, $J_{\downarrow}=1$ and $J_{\uparrow}=0.5$. The effective NN interactions $V_{\textrm{eff}}$ are marked in each panels, and the effective hopping amplitudes are all $J_{\textrm{eff}}=-0.125$. The doublons scattered bands and bound-doublon energy levels are indicated by the black and red lines, respectively. Here the energy is in units of $J_{\downarrow}$. (d) Parameter regions of hopping amplitudes where the systems support bound doublons BOs (gray area) and resonant doublons BOs (orange line). The markers ‘$\times$”, $\cdot$” and “$+$” indicate the regions of parameters for (a), (b) and (c), respectively. Here the unit is set as $\sqrt{UV}$, and strong on-site interaction and $U<0$, $V<0$ are assumed.
@@
$\beta(K)=\ln\left[\frac{V_{\mathrm{eff}}}{J_{\mathrm{eff}}\cos\left(K/2\right)}\right],$ (11)
@@
$E_{\textrm{bou.}}(K)=2V_{\mathrm{eff}}+\frac{2J_{\mathrm{eff}}^{2}\cos^{2}\left(K/2\right)}{V_{\mathrm{eff}}}.$ (12)
The bound-state condition $\beta(K)>0$ requires $\left|V_{\mathrm{eff}}\right|>\left|J_{\mathrm{eff}}\cos\left(K/2\right)\right|$. In this case, the doublons bound states are induced by the NN interaction $V$. In Fig. 2, we present the two-doublon spectrum for various NN interaction strengths, which can be obtained by diagonalizing the two-doublon effective Hamiltonian Eq. (7) or solving Eq. (10). In Fig. 2(a), we can see that when $V_{\mathrm{eff}}=0$, the spectrum is all composed of scattered band of doublons. Increasing $\left|V\right|$ would induce the isolated energy levels, which are the energies of bound doublons [see Fig. 2(b)]. When $\left|V_{\mathrm{eff}}\right|>\left|J_{\mathrm{eff}}\cos\left(K/2\right)\right|$ for all $K$, that is, $\left|V_{\mathrm{eff}}\right|>\left|J_{\mathrm{eff}}\right|$, the bound-doublon band is completely isolated from the scattered band [see Fig. 2(c)]. We illustrate these two parameter regions in Fig. 2(d) in units of $\sqrt{UV}$. In the following, we show that the systems in the two parameter regions support bound and resonant doublons BOs, respectively.
Here we first focus on the dynamics of the two bound doublons. For a large NN interaction $V$, we have $\left|V_{\mathrm{eff}}\right|\gg\left|J_{\mathrm{eff}}\cos\left(K/2\right)\right|$. In this case, two doublons in the bound band are strongly bound. Then we are able to write down the two-doublon effective Hamiltonian under the strongly bound basis $\left|\varphi_{j}\right\rangle=\hat{d}_{j}^{\dagger}\hat{d}_{j+1}^{\dagger}\left|\text{{Vac}}\right\rangle$. We divide the effective Hamiltonian Eq. (5) into two parts, i.e., $\hat{H}_{\mathrm{eff}}^{(0)}=V_{\mathrm{eff}}\sum_{j}\hat{n}_{j}\hat{n}_{j+1}/2$ and $\hat{H}_{\mathrm{eff}}^{\prime}=J_{\mathrm{eff}}\sum_{j}(\hat{d}_{j}^{\dagger}\hat{d}_{j+1}+\hat{d}_{j+1}^{\dagger}\hat{d}_{j})$, and treat $\hat{H}_{\mathrm{eff}}^{\prime}$ as a perturbation. According to the degenerate perturbation theory, the effective Hamiltonian is given by the following formula
@@
up to second order $J_{\mathrm{eff}}^{2}/V_{\mathrm{eff}}$. Here the projectors are defined as $\hat{P}=\sum_{j}\left|\varphi_{j}\right\rangle\left\langle\varphi_{j}\right|$ and $\hat{Q}=\hat{1}-\hat{P}$, and the energy of the unperturbed state is $E_{\mathrm{eff}}^{(0)}=2V_{\mathrm{eff}}$. Direct calculation shows that
$\hat{H}_{\mathrm{eff}}^{(2,\mathrm{bou.})}\approx\frac{J_{\mathrm{eff}}^{2}}{2V_{\mathrm{eff}}}\sum_{j}\left(\left|\varphi_{j}\right\rangle\left\langle\varphi_{j+1}\right|+\left|\varphi_{j+1}\right\rangle\left\langle\varphi_{j}\right|\right),$ (14)
in which the constant potential that is proportional to $\sum_{j}\left|\varphi_{j}\right\rangle\left\langle\varphi_{j}\right|$ is omitted. Now the bound-doublon physics is clear. From Eq. (14), we can see the bound-doublon hops in the lattice through a second-order process of the single-doublon hopping.
Figure 3: Numerical results of the time evolution of doublon density defined in Eq. (17) for different initial states with (a) delta distribution in Eq. (15) and (b) Gaussian distribution in Eq. (16). Parameters of the system are taken as $L=30$, $U=-6$, $V=-2$, $J_{\downarrow}=1$ and $J_{\uparrow}=0.5$. The effective NN interaction is $V_{\textrm{eff}}\approx-1.79$, and the effective hopping amplitude is $J_{\textrm{eff}}\approx-0.167$. The strength of tilted potential is taken as $F=0.001$. Parameters of the Gaussian wave packet are $\alpha=0.4$, $j_{0}=15$ and $K_{0}=\pi/2$. The time unit is $J_{\downarrow}^{-1}$, and we set $J_{\downarrow}=1$ throughout the numerical simulations in this work.
Figure 4: Numerical results of the time evolution of doublon density for different NN interaction $V$, tilted potential $F$ and doublon number. The system sizes for the two-doublon cases in (a)-(d) and four-doublon cases in (e) and (f) are taken as $L=30$ and $L=20$, respectively. Other system parameters are $U=-10$, $J_{\downarrow}=1$ and $J_{\uparrow}=0.1$. The effective NN interactions $V_{\textrm{eff}}$ are marked in each panels, and the effective hopping amplitudes are all $J_{\textrm{eff}}=-0.02$.
@@
$\left|\Psi_{\mathrm{a}}\left(0\right)\right>=\hat{d}_{j_{0}}^{\dagger}\hat{d}_{j_{0}+1}^{\dagger}\left|\text{{Vac}}\right\rangle,$ (15)
@@
Experimentally, the doublons density and its correlation can be directly measured in the Fermi-Hubbard lattice gas through the spin- and density-resolved quantum gas microscopy [84, 85, 86]. The numerical results are presented in Fig. 3, which are consistent with our analysis. Figures. 3(a) and (b) show that two different initial excitations undergo the breathing mode and oscillating mode, respectively. The oscillating periods are both $T=2\pi/(4F)\approx 1570.80$.
Figure 5: Mean-square displacement between the doublons [see Eq. (18)] as a function of time for the cases in Fig. 3(a) with parameters $U=-6$, $V=-2$, $J_{\downarrow}=1$, $J_{\uparrow}=0.5$, $F=0.001$, $V_{\textrm{eff}}\approx-1.79$ and $J_{\textrm{eff}}\approx-0.167$. and Figs. 4(a)-(d) with parameters $U=-10$, $J_{\downarrow}=1$, $J_{\uparrow}=0.1$ and $J_{\textrm{eff}}=-0.02$. While $V=-0.13$, $F=0$, $V_{\textrm{eff}}=-0.029$ for Fig. 4(a); $V=-0.13$, $F=0.005$, $V_{\textrm{eff}}=-0.029$ for Fig. 4(b); $V=-0.101$, $F=0$, $V_{\textrm{eff}}=0$ for Fig. 4(c); $V=-0.101$, $F=0.005$, $V_{\textrm{eff}}=0$ for Fig. 4(d). The system sizes are all $L=30$. The time scales are normalized by the final time of the data in each case.